OpenStudy (chrisplusian):

HELP PLEASE!!!!! Find the vectors T, N, and B at the given point... r(t) = 7cost, 7sint, 7lncost (7,0,0)

OpenStudy (chrisplusian):

For the unit tangent vector I got \[T(t)= <\cos^2 (t)-\sin^2(t), -2\cos(t)\sin(t),cost>\] Then evaluating for T(0) I got <0,1,0> and webassign said that answer was correct. I evaluated T(0) because I set the i-th component equal to the x value of the point. Then I set the j-th component equal to the y value of the point and both indicate that the value of t=0-degrees.